A company is planning on expanding and building a new plant in one of three Southeast Asian countries. Chris​ Ellis, the manager charged with making the​ decision, has determined that five critical success factors​ (CSFs) can be used to evaluate the prospective countries. Ellis used a rating system of 1​ (least desirable​ country) to 5​ (most desirable) to evaluate each CSF.

                                                                                                                                                                     

Candidate Country Ratings

Factor

Critical Success Factors

Weight

TaiwanTaiwan

ThailandThailand

SingaporeSingapore

1

Technology

0.15

5

4

3

2

Level of education

0.10

4

3

4

3

Political and legal aspects

0.40

4

3

4

4

Social and cultural aspects

0.10

5

4

3

5

Economic factors

0.25

2

3

1

​a) Based on the given​ information, the best location for Chris Ellis is

Thailand

Singapore

Taiwan

​, with a total weighted score of

nothing.

​(Enter your response rounded to two decimal​ places.)

​b) Political unrest in

TaiwanTaiwan

results in a lower​ score,

3​,

for Political and legal aspects. The best location for Chris Ellis is

Singapore

Taiwan

Thailand

​, with a total weighted score of

nothing.

​(Enter your response rounded to two decimal​ places.)

​c) If

TaiwanTaiwan​'s

score drops even​ further, to a​ 1, for Political and legal​ aspects, the best location for Chris Ellis is

Singapore

Taiwan

Thailand

​, with a total weighted score of

nothing.

​(Enter your response rounded to two decimal​ places.)
Answer #1

Solution:

Total weighted score is calculated as,

Total weighted score = Sum of (Weight x Country Ratings)

(a) Using the given information, the total weighted score of each location is calculated as below:

Taiwan = [(0.15 x 5) + (0.10 x 4) + (0.40 x 4) + (0.10 x 5) + (0.25 x 2)] = 3.75 (highest)

Thailand = [(0.15 x 4) + (0.10 x 3) + (0.40 x 3) + (0.10 x 4) + (0.25 x 3)] = 3.25

Singapore = [(0.15 x 3) + (0.10 x 4) + (0.40 x 4) + (0.10 x 3) + (0.25 x 1)] = 3.00

Based on the above calculations, the best location for Chris Ellis is Taiwan with a total weighted score of 3.75.

(b) Due to political unrest in Taiwan, its “Political and legal aspects” score drops to 3, the new total weighted score of each location is calculated as below:

Taiwan = [(0.15 x 5) + (0.10 x 4) + (0.40 x 3) + (0.10 x 5) + (0.25 x 2)] = 3.35 (highest)

Thailand = [(0.15 x 4) + (0.10 x 3) + (0.40 x 3) + (0.10 x 4) + (0.25 x 3)] = 3.25

Singapore = [(0.15 x 3) + (0.10 x 4) + (0.40 x 4) + (0.10 x 3) + (0.25 x 1)] = 3.00

Based on the above calculations, the best location for Chris Ellis is still Taiwan with a total weighted score of 3.35.

(c) If Taiwan’s “Political and legal aspects” score further drops to 1, the new total weighted score of each location is calculated as below:

Taiwan = [(0.15 x 5) + (0.10 x 4) + (0.40 x 1) + (0.10 x 5) + (0.25 x 2)] = 2.55

Thailand = [(0.15 x 4) + (0.10 x 3) + (0.40 x 3) + (0.10 x 4) + (0.25 x 3)] = 3.25 (highest)

Singapore = [(0.15 x 3) + (0.10 x 4) + (0.40 x 4) + (0.10 x 3) + (0.25 x 1)] = 3.00

Based on the above calculations, the best location for Chris Ellis is now Thailand with a total weighted score of 3.25.